OpenStudy (anonymous):

Can someone please help me? Given f(x) = x^2 + 4x - 1 and g(x) = 3x - 4, find (gog)(x)

geerky42 (geerky42):

\((g\circ g)(x) = g(~g(x)~) = g(~3x-4~) = 3(3x-4)-4\) Can you simplify \(3(3x-4)-4\)?

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

9x-16

geerky42 (geerky42):

Correct.

OpenStudy (anonymous):

mind helping me an other one?

OpenStudy (anonymous):

Given ƒ(x) = x2 + 4 and g(x) = 5x − 3, find (ƒ ○ ƒ)(x)